OpenStudy (anonymous):

Hi can someone help me determine how to solve this ? Determine an equation of a sphere given that one of the diameters has the extremities: (2,1,4) and (4,3,10)

zepdrix (zepdrix):

What does extremities mean? From one end to the other? Like the diameter or something? 0_o

OpenStudy (anonymous):

I'm not sure ... It's translated from french

OpenStudy (anonymous):

I have the answer though if you want

zepdrix (zepdrix):

Oh sure :)

OpenStudy (anonymous):

I just don't know how to get it loll

OpenStudy (anonymous):

\[(x-3)^{2}+(y -2)^{2}+(z-7)^2 = 11\]

zepdrix (zepdrix):

Oo nice I got the right answer :) Ok let's see if I can explain this in a way that makes sense.

OpenStudy (anonymous):

Nice.. All i know is finding the radius .. but after i'm not sure what to do

zepdrix (zepdrix):

So we'll start by finding the `diameter`. They gave us the coordinates from one end of the sphere to the other. Using the distance formula will give us the diameter.

OpenStudy (anonymous):

sqrt of 44?

zepdrix (zepdrix):

\[\Large\bf\sf (2,1,4),\qquad (4,3,10)\] \[\Large\bf\sf d\quad=\quad \sqrt{(4-2)^2+(3-1)^2+(10-4)^2}\]

zepdrix (zepdrix):

Yes good good. Let's simplify that value a lil bit.

zepdrix (zepdrix):

\[\Large\bf\sf \sqrt{44}\quad=\quad \sqrt{4\cdot11}\]

zepdrix (zepdrix):

We have a perfect square we can pull out of the root, yes?

OpenStudy (anonymous):

Agreed

zepdrix (zepdrix):

\[\Large\bf\sf diameter=2\sqrt{11}\]Ok good good. So our radius will be half that length.\[\Large\bf\sf radius=\sqrt{11}\]

zepdrix (zepdrix):

Then we need to find the location of the center of the sphere. We can do that using the midpoint formula thing.

zepdrix (zepdrix):

Just take the average of each coordinate.

OpenStudy (anonymous):

is that (x+y)/2

OpenStudy (anonymous):

Sorry x1+x2 / 2^

zepdrix (zepdrix):

Yes good, you've got the right idea. Average of each coordinate. \[\Large\bf\sf midpoint=\left(\frac{x_1+x_2}{2},\;\frac{y_1+y_2}{2},\;\frac{z_1+z_2}{2}\right)\]

OpenStudy (anonymous):

Thus giving (x-3)+(y-2)+(z-7)=11

OpenStudy (anonymous):

Awesome Thanks :D

zepdrix (zepdrix):

Yay team \c:/

OpenStudy (anonymous):

But the one thing i don<t get is why we need to write = 11 at the end

zepdrix (zepdrix):

\[\Large\bf\sf x^2+y^2+z^2=r^2\]We're shifting the center, so that's what the subtraction is about. See how our radius is being squared? That's why the square root disappears on the 11.

OpenStudy (anonymous):

Ohh I see .. basically it satisfies the general formula

zepdrix (zepdrix):

\[\Large\bf\sf r\quad=\quad \sqrt{11}\]\[\Large\bf\sf center\quad=\quad (3,2,7)\] Yes :) We just needed these two pieces of information so we could put it in standard form.

zepdrix (zepdrix):

I probably should have written out the standard form of a sphere with the h, k and l.. but whatever :P

OpenStudy (anonymous):

Loll It's perfect I get it :D

OpenStudy (anonymous):

Thanks a lot !!!!:D